OpenStudy (anonymous):

solve the absolute value equation... 2|4x-5| -2=4 explain how as well?

OpenStudy (anonymous):

1) add 2 to both sides 2) divide both sides by 2 3) solve \(4x-5=3\) and also \(4x-5=-3\) to finish

OpenStudy (anonymous):

how would i do |4x+1|=-3?